OpenStudy (goformit100):

In a cricket match, a batswoman hits a boundary 6 times out of 30 balls she plays. Find the probability that she did not hit a boundary.

OpenStudy (goformit100):

I am weak at probability please Help Me

OpenStudy (anonymous):

30 - 6 = 24 The probability that she did not hit the boundary is 24/30. Can also be written as 12/15 Hope this helps :)

OpenStudy (anonymous):

Probability of hitting outside the boundary based on given data is 6/30 or 1/5. So probability of not hitting outside the boundary would be 5/5 - 1/5 = 4/5

OpenStudy (goformit100):

Thank you sir.